# Optimization via Quantum Preconditioning

### v1

@@ B.1 Large-Depth Quantum Preconditioning at Small Scale
The focus of the main text was on large problems at shallow circuit depth $p$. We now turn our attention to quantum preconditioning of small-scale ($N=16$) random $3$-regular graph maximum cut problems and SK spin glasses for large values of $p$. We access large $p$ values via classical emulations using tabulated near-optimal QAOA angles up to $p\leq 11$ for random $3$-regular graph maximum cut problems [52, 29] and $p\leq 17$ for SK spin glasses [27, 28]. We consider simulated annealing as a solver.
We plot in Fig. 7 the average approximation ratio as a function of the number of iterations in simulated annealing for a fixed value of $p$ and as a function of $p$ for a fixed number of iterations. We observe that increasing $p$ in quantum preconditioning makes simulated annealing converges to an asymptotic approximation ratio value faster, i.e., in less iterations.
For a large number of iterations, we find that the approximation saturates to a value $\alpha<1$, as anticipated because original and preconditioned problems do not share the optimal solution except in the asymptotic $p\to+\infty$ limit. Nevertheless, we find that increasing $p$ increases the asymptotic approximation ratio value for SK spin glasses. It is less clear for the random $3$-regular graph maximum cut problems (Fig. 7a). We note that the random $3$-regular graph maximum cut problems are more likely to have degenerate solutions than the SK spin glasses beyond the global $\mathbb{Z}_{2}$ spin flip symmetry, due to the nature of their weights $\mathsf{W}_{ij}$: Unity versus normally distributed. As such, the correlation matrix $\mathsf{Z}^{(p)}$ may not tend to the desired asymptotic $p\to+\infty$ limit discussed in Sec. II: Another preconditioner may be more suited at large $p$.
@@ B.2 Solver-Agnostic Quantities
We consider solver agnostic quantities related to the problems hardness for random $3$-regular graph maximum cut problems and SK spin glasses. First, we compute the frustration index [161, 162, 163]
@@
where $\sigma_{i=1,2,\ldots N}\geq 0$ are the singular values sorted in ascending order of an input problem $\mathsf{I}+\mathsf{W}$ or $\mathsf{I}+\mathsf{Z}^{(p)}$. The normalized gap $\Delta\in[0,1]$ informs on how accurately the matrix can be approximated by a rank-$1$ matrix. Indeed, in the asymptotic $p\to+\infty$ limit in the presence of a doubly degenerate optimal solution ±$\pm\boldsymbol{z}_{\textbf{opt}}$, one has $\mathsf{I}+\mathsf{Z}^{(\infty)}=\boldsymbol{z}_{\textbf{opt}}\boldsymbol{z}_{\textbf{opt}}^{T}$, which is a rank-$1$ matrix, where in that case, the normalized gap is $\Delta=1$.
Figure 9: Average optimal solution overlap as defined in Eq. (14) as a function of the number of classical iterations in simulated annealing. (A) $N=128$-variable random $3$-regular graph maximum cut problems. (B) $N=128$-variable SK spin glasses. Each data point is averaged over $200$ randomly generated problem instances. Error bars indicate the standard error of the mean.
We plot in Fig. 8 these quantities on both the original problem $\mathsf{W}$ and preconditioned ones $\mathsf{Z}^{(p)}$. We find that for the largest $p$ values, the frustration index and gap asymptotically get to their $p\to+\infty$ value, i.e., $f\to 0$ and $\Delta\to 1$. This highlights the potential of quantum preconditioning to ease the hardness of challenging optimization problems. At a shallow circuit depth $p$, such as investigated in the rest of this work, the behavior of $f$ and $\Delta$ is much less clear. In particular, for random $3$-regular graph maximum cut problems, we observe an increase in frustration with $p$ at small $p$ in Fig. 8a: This is counterintuitive with the quantum preconditioned problem being actually easier to solve for algorithms such as simulated annealing (Fig. 1) and Burer-Monteiro (Fig. 2). However, we observe a change of slope for $p=2$, which indicates a drop in frustration for $p\geq 3$ and may result in a further ease of convergence at larger circuit depths for solvers relying on quantum preconditioning. Testing this hypothesis is out-of-scope for this work given the challenges of scaling up quantum circuits to $p=3$ and beyond on relevant large problem sizes $N$—see discussions in Sec. IV.2.
@@ B.3 Optimal Solution Overlap
Next, we consider the overlap of a sampled solution $\boldsymbol{z}$ via simulated annealing with the optimal one $\boldsymbol{z}_{\textbf{opt}}$. It is defined as
@@
reminiscent of the Edwards-Anderson overlap [164, 165, 166]. We plot in Fig. 9 the overlap with the optimal solution as a function of the number of classical iterations in simulated annealing for $N=128$ random $3$-regular graph maximum cut problems and SK spin glasses.
@@
In both cases, we find that for a fixed number of iterations of the classical solver, quantum preconditioning yields solutions $\boldsymbol{z}$ with a greater overlap to the optimal solution than simulated annealing working on the original problem. The overlap follows the trend of the approximation ratio in Figs. 13c and. 15c where we anticipate the original problem to asymptotically reach $q^{2}\to 1$ while the preconditioned problems, not sharing the same optimal solutions, will result in $q^{2}<1$. The larger overlap supports quantum preconditioning easing the convergence of classical solvers toward near-optimal solutions.
@@
Appendix C Experimental Implementation of the South Carolina Maximum Power Exchange Section Problem
@@
We implement quantum preconditioning at $p=1$ on Rigetti Ankaa-3 superconducting quantum computer for the South Carolina grid energy problem of Sec. III.3. Using the light-cone decomposition technique for computing necessary expectation values, this requires executing $600$ QAOA circuits ranging from $N=4$ to $N=14$ qubits (see Tab. 1). We use the transpilation strategy of Sec. IV.2.1 in terms of one- and two-qubit gates. The average fidelity of the two-qubit ISWAP gates is $\simeq 98.6\%$, of the one-qubit gate RX is $\simeq 99.8\%$, and of readout is $\simeq 96.7\%$ [167]. Computations are performed on linear chain of qubits with the phase separator of the QAOA compiled via an efficient swap network [79, 80]. Through a naïve swap network, $N$-qubit QAOA circuits at $p=1$ require $23N^{2}/$ 2 two-qubit ISWAP gates [70]—i.e., $24$ for $N=4$ and $294$ for $N=14$. Instead, the efficient swap network strategy makes this number range between $6$ and $83$ for the $600$ circuits at hand—this is a three to four fold improvement. We collect $K=10^{4}$ bit strings for each circuit for estimating expectation values.
@@
We show in Fig. 10a the experimental two-point correlation versus the exact one obtained via classical state vector emulations. As discussed in Sec. IV.2.2, we anticipate a simple depolarizing noise model to rescale expectation values by a factor corresponding to the circuit fidelity. We find that the data roughly follows a linear relationship with a slope $\mathcal{F}=0.718(7)$ obtained by least-squares fitting. In Fig. 10b, we rescale the exact value of each correlation by the corresponding circuit fidelity $F_{ij}$, which has been evaluated independently by multiplying the fidelity of all present individual operations, including one-qubit gates, two-qubit gates, and readout. The average fidelity over the $600$ circuits is $\mathbb{E}[F_{ij}]=0.679(6)$, which is relatively close to $\mathcal{F}$ and explains why the rescaled data of Fig. 10b follows more closely the $y=x$ line.
@@
Having preconditioned the problem, we now solve it using the classical BM algorithm—a comparison against SA in Fig. 4 showed that BM outperformed it by one order of magnitude, thus favoring BM as the solver of choice. We plot in Fig. 11 the approximation ratio versus the number of iterations. We find that the experimentally preconditioned problem provides an advantage when compared to the original problem. We observe that the preconditioned problem rescaled by the circuit fidelity does not yield an advantage here. At current noise levels, the benefit of preconditioning by implementing the QAOA in experiment is lower than that of the exact classical state vector emulation. However, we emphasize that we are achieving average approximation ratios above $\alpha\simeq 99.97\%$ and that the observed differences are extremely small.
@@
This highlights the experimental applicability of the proposed quantum preconditioning method, although more work is needed for quantum preconditioning to potentially deliver utility. Understanding how noise affects quantum preconditioning would be valuable.
Appendix D Implementation of Classical Solvers
@@
D.1 Simulated Annealing
Figure 11: Average approximation ratio against the number of iterations for the BM solver for the South Carolina grid energy problem. Different input problems are considered: Original problem, $p=1$ preconditioned problem via exact circuit emulations, $p=1$ preconditioned problem via experimental computations, $p=1$ preconditioned problem via exact circuit emulations with values rescaled by the circuit fidelity. Each data point is averaged over $100$ random independent run of the BM solver. Error bars indicate the standard error of the mean.
Simulated annealing is a physics-inspired solver treating an objective function such as Eq. (1) as a classical Ising model from which one wants to find the ground state. The search is based on a Metropolis-Hastings Markov chain Monte-Carlo algorithm [50, 51] with a Boltzmann distribution. It starts at a high temperature reduced throughout the iterations according to a predefined schedule such that one samples low-energy states, and ideally the ground state in the last step.
@@ D.1 Simulated Annealing
where $T_{\textrm{hot}}\geq T_{\textrm{cold}}\in\mathbb{R}$ are the initial and final temperatures, respectively. Finding optimal initial and final temperatures may require its own analysis and fine-tuning. The shape of the temperature schedule can also be optimized beyond a simple geometric descent. Here, we employ a heuristic to set $T_{\textrm{hot}}$ and $T_{\textrm{cold}}$ based on the $n$ edges and weights of the input problem. It has an algorithmic complexity of $O(n)$, as visible in Appendix H. This step, only performed once at the beginning, is independent of the number of sweeps and typically not a bottleneck of SA.
First, we define the effective field experienced by a spin $i$ due to two-body interactions with its nearest-neighbors
@@
In the context of the preconditioned problems one substitutes $\mathsf{W}_{ij}$ with $\mathsf{Z}_{ij}^{(p)}$ as defined in Eq. (2). Initially, when the temperature is hot, we want fast mixing such that all spins can be flipped with a probability $50\%$ (number chosen arbitrarily). Thus, $0.5=\exp(-2\max\nolimits_{i}h_{i}/T_{\textrm{hot}})$, where $2\max\nolimits_{i}h_{i}$ corresponds to the largest possible energy gap between two configurations after a spin flip. Inverting the relation, one gets
@@
At the coldest (and final) temperature, we bound the probability to excite any of the $N$ spins by $1\%$. Assuming that only spins with a minimal energy gap are excitable, one has $0.01=\nu\exp(-2\min\nolimits_{i}h_{i}/T_{\textrm{cold}})$, where we have approximated the minimal energy gap by $2\min\nolimits_{i}h_{i}$. Inverting the relation, one gets
$T_{\textrm{cold}}=2\min\nolimits_{i}h_{i}\bigr{/}\ln\bigl{(}100\nu\bigr{)},$ (18)
@@
where $\nu\in[1,N]$ counts the number of values $h_{i}$ equal to $\min_{i}h_{i}$, i.e., spins with a minimal energy gap. For instance, for the original random $3$-regular graph maximum-cut problems $h_{i}=3\forall i$ such that $T_{\textrm{hot}}=6\ln 2$ and $T_{\textrm{cold}}=6/\ln(100N)$ since $\nu=N$ in that case.
@@
The SA implementation used throughout this work relies on the Python package dwave-samplers 1.4.0 [168], although the underlying code is in C++ for efficiency. The temperature setting heuristic described above is the default one in this implementation.
D.2 Burer-Monteiro Solver
Figure 12: (A) Average number of nonzero terms in quantum preconditioned random $3$-regular graph maximum-cut problems for various values of $p$ as a function of the number of variables $N$. (B) Average number of terms to precondition in quantum preconditioned random $3$-regular graph maximum-cut problems for various values of $p$ as a function of the number of variables $N$. It is smaller or equal to the number of nonzero term via caching the preconditioned value of light-cone induced tree subgraphs. Each data point is averaged over $200$ randomly generated problem instances. Error bars indicate the standard error of the mean.

### v2

@@ B.1 Large-depth quantum preconditioning at small scale
The focus of the main text was on large problems at shallow circuit depth $p$. We now turn our attention to quantum preconditioning of small-scale ($N=16$) random $3$-regular graph maximum-cut problems and SK spin glasses for large values of $p$. We access large $p$ values via classical emulations using tabulated near-optimal QAOA angles up to $p\leq 11$ for random $3$-regular graph maximum-cut problems [51, 29] and $p\leq 17$ for SK spin glasses [27, 28]. We consider simulated annealing as a solver.
In Fig. 7, we plot the average approximation ratio as a function of the number of iterations in simulated annealing for a fixed value of $p$ and as a function of $p$ for a fixed number of iterations. We observe that increasing $p$ in quantum preconditioning makes simulated annealing converge to an asymptotic approximation ratio value faster, i.e., in fewer iterations.
For a large number of iterations, we find that the approximation saturates to a value $\alpha<1$, as anticipated, because the original and preconditioned problems do not share the optimal solution except in the asymptotic $p\to+\infty$ limit. Nevertheless, we find that increasing $p$ increases the asymptotic approximation ratio value for SK spin glasses. It is less clear for the random $3$-regular graph maximum-cut problems (Fig. 7a). We note that the random $3$-regular graph maximum-cut problems are more likely to have degenerate solutions than the SK spin glasses beyond the global $\mathbb{Z}_{2}$ spin flip symmetry, due to the nature of their weights $\mathsf{W}_{ij}$: unity versus normally distributed. As such, the correlation matrix $\mathsf{Z}^{(p)}$ may not tend to the desired asymptotic $p\to+\infty$ limit discussed in Sec. II. Another preconditioner may be more suited at large $p$.
@@ B.2 Solver-agnostic quantities
We consider solver-agnostic quantities related to the problem’s hardness for random $3$-regular graph maximum-cut problems and SK spin glasses. First, we compute the frustration index [159, 160, 161]
@@
where $\sigma_{i=1,2,\ldots N}\geq 0$ are the singular values sorted in ascending order of an input problem $\mathsf{I}+\mathsf{W}$ or $\mathsf{I}+\mathsf{Z}^{(p)}$. The normalized gap $\Delta\in[0,1]$ indicates how accurately the matrix can be approximated by a rank-$1$ matrix. Indeed, in the asymptotic $p\to+\infty$ limit in the presence of a doubly degenerate optimal solution $\pm\boldsymbol{z}_{\textbf{opt}}$, one has $\mathsf{I}+\mathsf{Z}^{(\infty)}=\boldsymbol{z}_{\textbf{opt}}\boldsymbol{z}_{\textbf{opt}}^{T}$, which is a rank-$1$ matrix, where in that case, the normalized gap is $\Delta=1$.
Figure 9: Average optimal solution overlap as defined in Eq. (14) as a function of the number of classical iterations in simulated annealing. (A) $N=128$-variable random $3$-regular graph maximum-cut problems. (B) $N=128$-variable SK spin glasses. Each data point is averaged over $200$ randomly generated problem instances. Error bars indicate the standard error of the mean.
In Fig. 8, we plot these quantities on both the original problem $\mathsf{W}$ and the preconditioned ones $\mathsf{Z}^{(p)}$. We find that for the largest $p$ values, the frustration index and gap asymptotically reach their $p\to+\infty$ value, i.e., $f\to 0$ and $\Delta\to 1$. This highlights the potential of quantum preconditioning to ease the hardness of challenging optimization problems. At a shallow circuit depth $p$, such as investigated in the rest of this work, the behavior of $f$ and $\Delta$ is much less clear. In particular, for random $3$-regular graph maximum-cut problems, we observe an increase in frustration with $p$ at small $p$ in Fig. 8a. This is counterintuitive, as the quantum-preconditioned problem is actually easier to solve for algorithms such as simulated annealing (Fig. 1) and Burer-Monteiro (Fig. 2; however, we observe a change of slope for $p=2$, which indicates a drop in frustration for $p\geq 3$, which may further facilitate convergence at larger circuit depths for solvers relying on quantum preconditioning. Testing this hypothesis is out of scope for this work given the challenges of scaling up quantum circuits to $p=3$ and beyond on relevant large problem sizes $N$ (see discussions in Sec. IV.2).
@@ B.3 Optimal Solution Overlap
Next, we consider the overlap of a sampled solution $\boldsymbol{z}$ via simulated annealing with the optimal one $\boldsymbol{z}_{\textbf{opt}}$. This is defined as
@@
reminiscent of the Edwards-Anderson overlap [162, 163, 164]. In Fig. 9, we plot the overlap with the optimal solution as a function of the number of classical iterations in simulated annealing for $N=128$ random $3$-regular graph maximum-cut problems and SK spin glasses.
@@
In both cases, we find that for a fixed number of iterations of the classical solver, quantum preconditioning yields solutions $\boldsymbol{z}$ with a greater overlap with the optimal solution than simulated annealing working on the original problem. The overlap follows the trend of the approximation ratio in Figs. 14c and. 17c, where we anticipate that the original problem will asymptotically reach $q^{2}\to 1$, while the preconditioned problems, not sharing the same optimal solutions, will result in $q^{2}<1$. The larger overlap supports quantum preconditioning easing the convergence of classical solvers toward near-optimal solutions.
Appendix C On the Light-Cone Decomposition Technique
On sparse problem instances, such as random $3$-regular graphs and the South Carolina grid network, one can leverage a light-cone decomposition technique for estimating expectation values such as $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle$ for preconditioning the problem.
Indeed, a $p$-layer QAOA circuit as defined in Eq. (3) induces quantum coherence around each qubit and its $p$-nearest neighbors, as per the underlying graph topology defined by $\mathcal{G}$. This defines a light-cone-induced subgraph $\mathcal{G}_{i}$ for each of the qubits or problem variables $i$, which can straightforwardly be obtained. For computing an expectation value $\langle\hat{Z}_{i}\hat{Z}_{j}\cdots\hat{Z}_{m}\rangle$, one may only consider the union of the light-cone-induced subgraphs for each of the involved variables, i.e., $\mathcal{G}_{ij\cdots m}=\mathcal{G}_{i}\cup\mathcal{G}_{j}\cdots\cup\mathcal{G}_{m}$. Executing the QAOA on either $\mathcal{G}$ or $\mathcal{G}_{ij\cdots m}$ will lead to the same expectation value $\langle\hat{Z}_{i}\hat{Z}_{j}\cdots\hat{Z}_{m}\rangle$.
The light-cone technique presents an advantage for computing $\langle\hat{Z}_{i}\hat{Z}_{j}\cdots\hat{Z}_{m}\rangle$ when $\mathcal{G}_{ij\cdots m}$ is smaller than $\mathcal{G}$, as the resulting QAOA execution involves fewer qubits and gates. The price is that of a different QAOA execution for every single desired expectation value $\langle\hat{Z}_{i}\hat{Z}_{j}\cdots\hat{Z}_{m}\rangle$ corresponding to different subgraphs $\mathcal{G}_{ij\cdots m}$.
For a graph with $N$ nodes with a maximum degree $d$, the number of nodes in a light-cone-induced subgraph by the QAOA with depth $p$ for computing a two-point correlation $\langle\hat{Z}_{i}\hat{Z}_{j}\rangle$ is at most
$\textrm{min}\left(1+\frac{2d\left[(d-1)^{p}-1\right]}{d-2},N\right).$ (15)
For example, this upper bound is saturated for $d$-regular graphs such as the random $3$-regular graph considered. For $N$-variable all-to-all graphs where d=N1d=N-1, such as the SK problems considered, the light-cone subgraphs involve all the $N$ nodes, meaning that the execution of the QAOA on the full graph is required. This worst-case bound is not saturated for the grid energy network. The light-cone technique is exact and leads to smaller simulations for sparse graphs at shallow QAOA depth $p$ while mapping to the standard QAOA for dense graphs or large $p$ values.
@@ Appendix C On the Light-Cone Decomposition Technique
Appendix D Experimental Implementation of the South Carolina Maximum Power Exchange Section Problem
@@
We implement quantum preconditioning at $p=1$ on the Rigetti Ankaa-3 superconducting quantum computer for the South Carolina grid energy problem of Sec. III.3. Using the light-cone decomposition technique for computing necessary expectation values, this requires executing $600$ QAOA circuits ranging from $N=4$ to $N=14$ qubits (see Tab. 1). We use the transpilation strategy of Sec. IV.2.1 in terms of one- and two-qubit gates. The average fidelity of the two-qubit ISWAP gates is approximately $98.6\%$, of the one-qubit gate RX is approximately $99.8\%$, and that of the readout is approximately $96.7\%$ [165]. Computations are performed on a linear chain of qubits with the phase separator of the QAOA compiled via an efficient swap network [78, 79]. Through a naïve swap network, $N$-qubit QAOA circuits at $p=1$ require $3N^{2}/2$ two-qubit ISWAP gates [69]—i.e., $24$ for $N=4$ and $294$ for $N=14$. Instead, the efficient swap network strategy makes this number range between $6$ and $83$ for the $600$ circuits at hand—this is a three-to-four-fold improvement. We collect $K=10^{4}$ bit strings for each circuit for estimating expectation values.
@@
In Fig. 10a, we show the experimental two-point correlation versus the exact one obtained via classical state-vector emulations. As discussed in Sec. IV.2.2, we anticipate a simple depolarizing noise model to rescale expectation values by a factor corresponding to the circuit fidelity. We find that the data roughly follows a linear relationship with a slope $\mathcal{F}=0.718(7)$ obtained by least-squares fitting. In Fig. 10b, we rescale the exact value of each correlation by the corresponding circuit fidelity $F_{ij}$, which has been evaluated independently by multiplying the fidelity of all present individual operations, including one-qubit gates, two-qubit gates, and readout. The average fidelity over the $600$ circuits is $\mathbb{E}[F_{ij}]=0.679(6)$, which is relatively close to $\mathcal{F}$ and explains why the rescaled data of Fig. 10b follows the $y=x$ line more closely.
@@
Having preconditioned the problem, we now solve it using the classical BM algorithm—a comparison against SA in Fig. 4 showed that BM outperformed it by one order of magnitude, thus favoring BM as the solver of choice. In Fig. 11, we plot the approximation ratio versus the number of iterations. We find that the experimentally preconditioned problem provides an advantage when compared to the original problem. We observe that the preconditioned problem rescaled by the circuit fidelity does not yield an advantage here. At current noise levels, the benefit of preconditioning by implementing the QAOA in experiment is lower than that of the exact classical state-vector emulation; however, we emphasize that we are achieving average approximation ratios above $\alpha\simeq 99.97\%$ and that the observed differences are extremely small.
@@
This highlights the experimental applicability of the proposed quantum-preconditioning method, although more work is needed for quantum preconditioning to potentially deliver utility. Understanding how noise affects quantum preconditioning would be valuable.
Appendix E Implementation of Classical Solvers
@@
E.1 Simulated annealing
Figure 11: Average approximation ratio against the number of iterations for the BM solver for the South Carolina grid energy problem. Different input problems are considered: the original problem, $p=1$ preconditioned problem via exact circuit emulations; the $p=1$ preconditioned problem via experimental computations; and the $p=1$ preconditioned problem via exact circuit emulations with values rescaled by the circuit fidelity. Each data point is averaged over $100$ random independent runs of the BM solver. Error bars indicate the standard error of the mean.
Simulated annealing is a physics-inspired approach to solving that treats an objective function such as Eq. (1) as a classical Ising model for which one wants to find the ground state. The search is based on a Metropolis-Hastings Markov-chain Monte Carlo algorithm [49, 50] with a Boltzmann distribution. It starts at a high temperature, which is reduced throughout the iterations according to a predefined schedule such that one samples low-energy states, and ideally the ground state in the last step.
@@ E.1 Simulated annealing
where $T_{\textrm{hot}}\geq T_{\textrm{cold}}\in\mathbb{R}$ are the initial and final temperatures, respectively. Finding optimal initial and final temperatures may require its own analysis and fine-tuning. The shape of the temperature schedule can also be optimized beyond a simple geometric descent. Here, we employ a heuristic to set $T_{\textrm{hot}}$ and $T_{\textrm{cold}}$ based on the $n$ edges and weights of the input problem. This has an algorithmic complexity of $O(n)$, as visible in Appendix K. This step, which is only performed once at the beginning, is independent of the number of sweeps and typically not a bottleneck for SA.
First, we define the effective field experienced by a spin $i$ due to two-body interactions with its nearest neighbors
@@
In the context of the preconditioned problems, one substitutes $\mathsf{W}_{ij}$ with $\mathsf{Z}_{ij}^{(p)}$ as defined in Eq. (2). Initially, when the temperature is hot, we want fast mixing such that all spins can be flipped with a probability of $50\%$ (number chosen arbitrarily). Thus, $0.5=\exp(-2\max\nolimits_{i}h_{i}/T_{\textrm{hot}})$, where $2\max\nolimits_{i}h_{i}$ corresponds to the largest possible energy gap between two configurations after a spin flip. Inverting the relation, one gets
@@
At the coldest (and final) temperature, we bound the probability of exciting any of the $N$ spins by $1\%$. Assuming that only spins with a minimal energy gap are excitable, one has $0.01=\nu\exp(-2\min\nolimits_{ij}|\mathsf{W}_{ij}|/T_{\textrm{cold}})$, where we have approximated the minimal energy gap by $2\min\nolimits_{ij}|\mathsf{W}_{ij}|$. Inverting the relation, one gets
$T_{\textrm{cold}}=2\min\nolimits_{ij}|\mathsf{W}_{ij}|\bigr/\ln\bigl(100\nu\bigr),$ (19)
@@
where $\nu\in[1,N]$, counts for $i=1,\ldots,N$, the number of values $\min_{i}|\mathsf{W}_{ij}|$ equal to $\min\nolimits_{ij}|\mathsf{W}_{ij}|$, i.e., spins with a minimal energy gap. For instance, for the original unit-weight random $3$-regular graph maximum-cut problems, $h_{i}=3\forall i$, such that $T_{\textrm{hot}}=6\ln 2$ and $T_{\textrm{cold}}=6/\ln(100N)$, since $\nu=N$ in this case.
@@
The SA implementation used throughout this work relies on the Python package dwave-samplers 1.4.0 [166], although the underlying code is in C++ for efficiency. The temperature-setting heuristic described above is the default one in this implementation.
Figure 12: Effect on the approximation ratio returned by simulated annealing versus the number of iterations due to variations in QAOA angles defined by $\varepsilon>0$, relative to the near-optimal ($\varepsilon=0$) case. SK problem instances with $N=64$ variables at $p=1$ are considered, where each data point is averaged over $500$ randomly generated instances. Error bars indicate the standard error of the mean.
E.2 Burer-Monteiro solver